what is the slope-intercept form of the line passing through the point (6,-9) and parallel to the line 5x-7y=-1
Answer by rmromero(383)


what is the slope-intercept form of the line passing through the point (6,-9) and parallel to the line 5x-7y=-1

Slope - Intercept Form
y = mx + b , where m = slope and b = y-intercept

We can find the slope of the line passing through (6, -9) and parallel
to the line 5x - 7y = -1, by finding the slope of the line parallel to it.
That is 5x - 7y = -1.
Find the slope of 5x - 7y = -1
5x + 1 = 7y
5 1
___x + ____ = y
7 7
Slope or m = 5/7
Since the two line are parallel, they have the same slope
so the slope of the line passing through (6,-9) and parallel to
the line 5x - 7y = -1 is 5/7.

Now, we will find the b, the y - intercept.
y = mx + b, where m= 5/7 and x = 6 and y = -9
-9 = 5/7(6) + b
-9 = 30/7 + b
-9 - 30/7 = b
-72/7 = b
The slope - intercept form of the new line is
y = (5/7)x - 72/7